OpenStudy (anonymous):

What is the quotient of 3 1/10 divided by 2/5 in simplest form? 7 3/4 7 1/2 3 1/4 3 1/25

OpenStudy (anonymous):

10 * 5/2 = 25 3/1*25/10 = 15/7 or 7.5